# Phonon Induced Contrast in Matter Wave Interferometer

### v1

@@ Abstract
Utilizing the Stern-Gerlach apparatus to create matter-wave superposition states is a long-sought-after goal, not only due to its potential applications in the quantum realm but also because of its fundamental implications for studying the quantum properties of gravity. The main challenge in creating a macroscopic quantum interferometer arises from the loss of coherence, primarily through two channels. One channel involves strong coupling with the environment for macroscopic matter, leading to decoherence. The other channel relates to the precision of wave packet overlap, which can occur due to external and internal fluctuations of various sources. The latter introduces a unique challenge for larger-scale masses by perturbing the centre of mass motion of the macroscopic object. Here, we study a particular challenge, namely, the issue of internal degrees of freedom, specifically phonon fluctuations and contrast reduction. This work will investigate the contrast reduction caused by spin-magnetic field and diamagnetic interactions at the phonon occupation level in the quantum gravity-induced entanglement of masses (QGEM) protocol configuration.
@@ I Introduction
The matter-wave interferometer is an essential tool for generating sizeable spatial superposition states. It represents the forefront of theoretical physics, as the creation of spatial superposition states can aid in constructing a theory that reconciles gravity with quantum mechanicsPenrose (1996); Diósi (1988); Pearle (1989); Bassi et al. (2013); Nimmrichter and Hornberger (2013), studying the equivalence principleOverstreet et al. (2018); Asenbaum et al. (2020); Bose et al. (2023); Chakraborty et al. (2023), decoherence mechanismRomero-Isart (2011); Romero-Isart et al. (2011, 2010); Tilly et al. (2021); Schut et al. (2022a); van de Kamp et al. (2020); Rijavec et al. (2021); Gunnink et al. (2023); Schut et al. (2022b, 2023a); Fragolino et al. (2023); Toroš et al. (2020); Schut et al. (2023b). In terms of applications, matter-wave interferometers are also considered for detecting gravitational wavesMarshman et al. (2020a), neutrinos Kilian et al. (2023), detecting light axion-like particles Barker et al. (2022), for precise measurements of Earth’s gravitational accelerationPeters et al. (1999); Chiao and Speliotopoulos (2004); Roura et al. (2006); Foffa et al. (2006); Dimopoulos et al. (2008, 2009); Tino et al. (2019); Graham et al. (2013); Marshman et al. (2020a), and detection of space debris Wu et al. (2023). To test the quantum nature of gravity, a proposal known as quantum gravity induced entanglement of masses (QGEM) has been recently proposed Bose et al. (2017)  [^1] This work was first reported in Bose (2016). A similar proposal was made in Ref. Marletto and Vedral (2017).. To further probe the nature of gravity, the gravitational-optomechanical test, which tests the analogue of light-bending experiment in quantum gravity Biswas et al. (2023), and measurement-based tests Hanif et al. (2023), all require the creation of a matter-wave interferometer. The QGEM proposal utilizes the two masses, each prepared in a spatial superposition state, which is kept at a distance where the only interaction between them is mediated via gravity. If detectable entanglement exists between these two masses, one can ascertain that the gravitational interaction is quantum in nature Marshman et al. (2020b); Bose et al. (2022); Vinckers et al. (2023); Elahi and Mazumdar (2023); Christodoulou et al. (2022); Carney et al. (2019); Danielson et al. (2022). This quantum-entanglement-based, experimental is concordant with the LOCC theorem, which states that classically one cannot mediate quantum entanglement between the two quantum systems Bennett et al. (1996). In QGEM, the spatial superposition is created by the Stern-Gerlach (SG) apparatus Friedrich and Schmidt-Böcking (2021) while the test mass is played by a diamond internally embedded with a nitrogen-vacancy centre (NV). The SG device utilizes the coupling between the magnetic field and the spin (NV-centre) to create a spatial superposition state. Based on similar mechanisms, there have been approaches toward atom interferometry. Such interferometer has been experimentally realized on atom chips for half-loop Machluf et al. (2013) and full-loop Margalit et al. (2020) schemes. For the former, the coherent time for maintaining spatial superposition was $21.45$ ms, and the size of 3.98 $\mu$m was reached. For the full loop, the coherent time and superposition size were 7 ms and 0.38 $mu$m, respectively. Of course, QGEM imposes higher demands on the interferometer’s performance, with critical requirements including the diamond’s mass $M\sim 10^{-15}-10^{-14}$ kg and the spatial superposition size needing to reach approximately $10-100\mu$Bose et al. (2017); van de Kamp et al. (2020); Schut et al. (2023b, 2022b), and Schut et al. (2023c). The larger mass objects tend to couple more strongly with their environment, such as the electromagnetic fields and collisions with residual gas particles. These couplings act as measurements, causing the collapse of the spatial superposition state into classical states, known as decoherence Joos (2003); Schlosshauer (2019). Furthermore, due to experimental constraints on the magnetic field gradient Marshman et al. (2022, 2023), the time required to achieve large-scale spatial superposition states is on the order of $\Delta t\sim 1$ second makes it more challenging to maintain the spin coherence state of NV centres and the spatial superposition state of diamond  [^2] See also gravity gradient and relative acceleration noise Toroš et al. (2021); Großardt (2020); Wu et al. (2023) for other sources of noise that lead to dephasing the interferometer along with electromagnetic sources of noise Schut et al. (2023a); Fragolino et al. (2023)..
@@
For a closed-loop scheme, another challenge that is independent of the environment arises from the Humpty-Dumpty (HD) effectEnglert et al. (1988); Schwinger et al. (1988); Scully et al. (1989); Englert (1997), which refers to the overlap problem of matter waves. For example, in the atomic interferometer mentioned above, if the wave packets on both sides of the interferometer arm cannot precisely overlap (both in translational and rotational degrees of freedom), they cannot generate interference fringes and thus lead to the loss of contrast, see Japha and Folman (2022) where the analysis of HD effect has been analysed quite nicely including the impact of angular momentum and the libration mode of the NV spin. For macroscopic objects, there is also a challenge of overlap due to the internal degrees of freedom, such as phonon vibration. In a very interesting workHenkel and Folman (2022), the contrast reduction of phonon wave packets induced by coupling NV centres with magnetic fields was demonstrated. More recently, this coupling has been extended to white noise, providing a boundary on creating larger-scale spatial superposition states that have been considered Henkel and Folman (2023). This work will also investigate the contrast loss caused by phonons in matter-wave interferometers. Unlike atomic-scale interferometers, due to the diamagnetic property of a diamond and its interaction with the magnetic field being proportional to its mass, the diamagnetic interaction becomes essential in the matter-wave interferometers with a massive object. For example, the diamagnetic interaction contributes to the trajectories of the diamond’s centre of mass (CoM) wave packets. Therefore, the magnetic field must be carefully designed to guarantee a close loopMarshman et al. (2022); Zhou et al. (2022a, b, 2023). Besides, there has been recently proposed a mass-independent schemeZhou et al. (2023) based on diamagnetic interaction of creating large superposition size. From the perspective of internal degrees of freedom, the diamagnetic repulsions of the diamonds on either side of the interferometer arms are different, leading to the excitation of varying phonon modes and thus resulting in a contrast reduction of internal degrees of freedom. This is fundamentally similar to the spin-magnetic field interaction papers Henkel and Folman (2022, 2023). In this regard, our paper differs from the Refs.Henkel and Folman (2022, 2023). We validate their results for the spin-magnetic field interaction. However, we also Consider the effect of diamagnetic contribution, which brings another layer of complexity that is not present otherwise. In the case of diamagnetic contribution, the phonons are also coupled to the CoM of the diamond, intrinsically bringing a new kind of interaction. The difference lies in the fact that for diamagnetic interaction, each carbon atom in the atomic chain on both sides of the interferometer arms experiences repulsion in the same direction, while for the spin-magnetic interaction, the chains are plucked at the NV-centre in opposite directions, as illustrated in Fig. 1. Besides the magnetic effects, we will also briefly mention the role of electric dipole interaction with phonons and how it affects the latter in comparing the loss of contrast for the one-loop interferometer. The structure of this work is arranged as follows. In section II, we introduced the setup, which includes the phonon model and the statistical model of phonon modes at finite temperature, and introduced the Wigner function. In III, we provided the Hamiltonian of the diamond and presented the model for the trajectory of the diamond wavepacket. In IV, we examined the different effects of diamagnetic repulsion and spin-magnetic field interaction on phonon dynamics and further utilized displacement operators in Section V to describe the phonon’s contrast reduction. In Section VI, we presented numerical results and compared the contrast loss induced by these two interactions at different temperatures. Also, we have considered contrast loss led by other interactions, potential dipole-electric field interaction, in the level of phonon; this part of the discussion can be found in Appendix C and D.
[…]
@@
Figure 1: Illustrations of different forces acting on a lattice chain (represented by black solid wave lines) where positions of atom/spin are dotted in black. The wave packets of the centre of mass (CoMs) in the two interferometry arms denoted here by the left (L) and right (R) arms are plotted by orange and blue circles. Panel (a) shows splitting forces $F_{s}^{\text{L,R}}$ acting on the spin with site number $s$ have the same magnitude but opposite direction. Panel (b) shows diamagnetic forces $F_{i}^{\text{L,R}}$ acting on every atom, where, for simplicity, we assume it acts on the spin as well without changing the main dynamics of the chain.
@@ II Brief review on Phonon
where $K_{ij}$ is a matrix associated with the inter-atomic interaction while $Q_{q}(X_{i})=Q_{q}^{i}$ represents the eigenvector of the matrix. The position of each atom can be represented as $X_{i}(t)+x_{i}(t)$, where the capital $X_{i}(t)$ denotes the equilibrium position of the $i^{\text{th}}$ atom at any instant of time, and $x_{i}$ represents its small vibrations. Because the position of the diamond as a whole varies within the SG apparatus, the equilibrium position $X_{i}(t)$ of each atom in the chain becomes a function of time. Notably, the relative distances between these equilibrium positions, $X_{i+1}-X_{i}=\text{const.}$., remain constant. Therefore, $X_{i}(t)$ can effectively be regarded as the trajectory of the diamond, with their distances relative to the centroid $X_{\text{c(enter)}}$ remaining unchanged. Here, for convenience, we assume that the NV centre within the atom is embedded at the centroid’s position, $X_{\text{s(pin)}}(t)=X_{\text{c}}(t)$. In the following, we sometimes neglect the time $t$ for convenience. The small movement $x_{i}(t)$ can be expanded into mode amplitude $u_{q}$, with creation and annihilation operators as,
@@
For simplicity, in our one-dimensional atomic chain model embedded with an NV centre, we assume that all particles have identical mass, and the total mass is given by $M=Nm$, where $N$ is the total number of particles in the diamond, and $m$ is the atomic mass of the diamond. The $\omega_{q}$ is the eigen-frequency of the chain model with fundamental tone $\omega_{0}={\pi c}/{L}$, where $c=17.5\times 10^{3}{\rm m/s}$ is the sound speed within a diamond Jishi, Radi A (2013) and $L$ is the length. For example, for the masses in a range $M\sim 10^{-14}\sim 10^{-20}$ kg the corresponding fundamental tone ranges from $\omega_{0}\sim 10^{10}-10^{12}$ Hz. At the initial stage, $t=0$, before the diamond enters the SG apparatus and interacts with the magnetic field, one can assume that the phonons are in a thermal equilibrium state at a temperature $T$ Jishi, Radi A (2013),
@@
where $\sigma_{u_{q}}$ and $\sigma_{\dot{u}_{q}}$ are statistical averages. A review of this detail can be found in the appendix A. In the case of thermal equilibrium, the phonons in each mode $q$ form a mixed state. The amplitudes and momenta of the phonons thus take on Gaussian distributions, represented by the Wigner functionCase (2008) as follows,
$W\left(u_{q},\dot{u}_{q}\right)={\cal{N}}\exp\left[-\left(\frac{u_{q}^{2}}{2\sigma_{u_{q}}^{2}}+\frac{\dot{u}_{q}^{2}}{2\sigma_{\dot{u}_{q}}}\right)\right],$ (6)
where $\cal{N}$ is normalization factor. Eq. (5) gives the normal mode’s characteristic length and momentum. Note that $\sigma_{u_{q}}$ and $\sigma_{\dot{u}_{q}}$ are statistical averages dependent on temperature $T$, see appendix A. Once we let the diamond experience the SG force, the occupation number of the phonons will evolve, and we will estimate that in the following sections.
@@ III Interactions in Stern-Gerlach Apparatus
For the diamond embedded with a single spin $S$ in the NV, we already assume that it has been prepared in a superposition of opposite direction $S^{\text{R}}=1$ (right) and $S^{\text{L}}=-1$ (left), hence the spatial superposition. The Hamiltonian for the diamond’s CoM consists of three parts  [^3] There is also a term related to $D(S\cdot\hat{n})^{2}$ related to the zero-point splitting Japha and Folman (2022); Zhou et al. (2022a). This term is relevant for the internal spin degrees of freedom. Here, we are not considering this contribution; we do not consider the rotation of the diamond in our current context.,
@@
the first term is just the kinetic energy relating to CoMs’s momentum. The second term is the spin-magnetic field coupling with the magnetic moment $\mu=g\mu_{\text{b}}$, where the Landé g-factor $g\approx 2$ and Bohr magneton $\mu_{\text{b}}=9.27\times 10^{-24}$ (J/T). For simplicity, we assume the NV-centre is located at the centre of mass of the diamond with equilibrium position $X_{s}=X_{c}$. The third term relates to diamagnetic energy with mass susceptibility $\chi_{\rho}=-6.2\times 10^{-9}{\rm m^{3}/kg}$ and vacuum permeability $\mu_{0}$. For the magnetic field in Eq. (7), we give its form that is experienced by the $i^{\text{th}}$ atom as,
@@
where $B_{0}$ is the bias magnetic field, and $b(t)X_{i}(t)$ denotes the gradient in the SG setup experienced by the $i^{\rm th}$ atom. Note that in Eq.(8), the coordinates along the $\mathbf{e}_{x}$-axis are represented by the uppercase $X_{i}$, indicating the equilibrium positions of atoms in the chain. For example, the magnetic field experienced by the spin, which we have assumed to be located at the centre of the chain, becomes $B_{\text{c}}=B_{0}+b(t)X_{\text{c}}$. In this model, we only consider the one-dimensional case, where the separation of the diamond and the force on the atomic chain is confined to the $\mathbf{e}_{x}$ direction. We do not consider the motion of the diamond in other dimensions, such as in drop tower experiments where the diamond needs to move in the direction of Earth’s acceleration or in magnetic levitation setups where it moves along a direction orthogonal to $\mathbf{e}_{x}$. In these cases, the form of a magnetic field becomes very complicated for a closed loop, and it needs a careful design, especially when one takes the diamagnetic term into account Zhou et al. (2022c, 2023). For simplicity, we consider the movements of CoMs only along $\mathbf{e}_{x}$, and they are subjected only to the spin-magnetism interaction (second term in Eq. (7)). Therefore, to get a closed loop, the magnetic gradient $b(t)$ can be expressed in a very simple form as Toroš et al. (2021),

### v2

# Phonon-induced contrast in a matter-wave interferometer

@@ Abstract
Utilizing the Stern-Gerlach apparatus to create matter-wave superposition states is a long-sought-after goal, not only due to its potential applications in the quantum realm but also because of its fundamental implications for studying the quantum properties of gravity. The main challenge in creating a macroscopic quantum interferometer arises from the loss of coherence, primarily through two channels. One channel involves strong coupling with the environment for macroscopic matter, leading to decoherence. The other channel relates to the precision of wave packet overlap, which can occur due to external and internal fluctuations of various sources. The latter introduces a unique challenge for larger-scale masses by perturbing the centre of mass motion of the macroscopic object. Here, we study a particular challenge, namely, the issue of internal degrees of freedom, specifically phonon fluctuations and contrast reduction. This work will investigate the contrast reduction caused by spin-magnetic field-phonon and diamagnetism-phonon interactions in the quantum gravity-induced entanglement of masses (QGEM) protocol configuration.
@@ I Introduction
The matter-wave interferometer is an essential tool for generating sizeable spatial superposition states. It represents the forefront of theoretical physics, as the creation of spatial superposition states can aid in constructing a theory that reconciles gravity with quantum mechanicsPenrose (1996); Diósi (1989); Pearle (1989); Bassi et al. (2013); Nimmrichter and Hornberger (2013), studying the equivalence principleOverstreet et al. (2018); Asenbaum et al. (2020); Bose et al. (2023); Chakraborty et al. (2023), decoherence mechanismRomero-Isart (2011); Romero-Isart et al. (2011, 2010); Tilly et al. (2021); Schut et al. (2022a); van de Kamp et al. (2020); Rijavec et al. (2021); Gunnink et al. (2023); Schut et al. (2022b, 2023a); Fragolino et al. (2023); Toroš et al. (2020); Schut et al. (2023b). In terms of applications, matter-wave interferometers are also considered for detecting gravitational wavesMarshman et al. (2020a), neutrinos Kilian et al. (2023), detecting light axion-like particles Barker et al. (2022), for precise measurements of Earth’s gravitational accelerationPeters et al. (1999); Chiao and Speliotopoulos (2004); Roura et al. (2006); Foffa et al. (2006); Dimopoulos et al. (2008, 2009); Tino et al. (2019); Graham et al. (2013); Marshman et al. (2020a), and detection of space debris Wu et al. (2023).
To test the quantum nature of gravity, a proposal known as quantum gravity induced entanglement of masses (QGEM) has been recently proposed Bose et al. (2017); Marletto and Vedral (2017). To further probe the nature of gravity, the gravitational-optomechanical test, which tests the analogue of light-bending experiment in quantum gravity Biswas et al. (2023), and measurement-based tests Hanif et al. (2023), all require the creation of a matter-wave interferometer.
The QGEM proposal utilizes the two masses, each prepared in a spatial superposition state, which is kept at a distance where the only interaction between them is mediated via gravity. Suppose detectable entanglement exists between these two masses. In that case, one can ascertain that the gravitational interaction is quantum in nature Marshman et al. (2020b); Bose et al. (2022); Vinckers et al. (2023); Elahi and Mazumdar (2023); Christodoulou et al. (2022); Carney et al. (2019); Danielson et al. (2022). This quantum-entanglement-based, experimental is concordant with the local operations and classical communication (LOCC) theorem, which states that classically one cannot mediate quantum entanglement between the two quantum systems Bennett et al. (1996).
In QGEM, the spatial superposition is created by the Stern-Gerlach (SG) apparatus Friedrich and Schmidt-Böcking (2021). At the same time, the test mass is played by a diamond internally embedded with a nitrogen-vacancy centre (NV). The SG device utilizes the coupling between the magnetic field and the spin (NV-centre) to create a spatial superposition state. Based on similar mechanisms, there have been approaches toward atom interferometry. Such interferometer has been experimentally realized on atom chips for half-loop Machluf et al. (2013) and full-loop Margalit et al. (2020) schemes. For the former, the coherent time for maintaining spatial superposition was $21.45$ ms, and the size of 3.98 $\mu$m was reached. For the full loop, the coherent time and superposition size were 7 ms and 0.38 $\mu$m, respectively.
Of course, QGEM imposes higher demands on the interferometer’s performance, with critical requirements including the diamond’s mass $M\sim 10^{-15}-10^{-14}$ kg (with length scale $\sim 10^{-7}-10^{-6}$ m) and the spatial superposition size needing to reach approximately $10-100\mu$Bose et al. (2017); van de Kamp et al. (2020); Schut et al. (2023b, 2022a), and Schut et al. (2023c). The larger mass objects tend to couple more strongly with their environment, such as the electromagnetic fields and collisions with residual gas particles. These couplings act as measurements, causing the collapse of the spatial superposition state into classical states, known as decoherence Joos (2003); Schlosshauer (2019). Furthermore, due to experimental constraints on the magnetic field gradient Marshman et al. (2022, 2023), the time required to achieve large-scale spatial superposition states is on the order of $\Delta t\sim 1$ second makes it more challenging to maintain the spin coherence state of NV centres and the spatial superposition state of diamond  [^1] See also gravity gradient and relative acceleration noise Toroš et al. (2021); Großardt (2020); Wu et al. (2023) for other sources of noise that lead to dephasing the interferometer along with electromagnetic sources of noise Schut et al. (2023a); Fragolino et al. (2023)..
@@
For a closed-loop scheme, another challenge that is independent of the environment arises from the Humpty-Dumpty (HD) effectEnglert et al. (1988); Schwinger et al. (1988); Scully et al. (1989); Englert (1997), which refers to the overlap problem of matter waves. For example, in the atomic interferometer mentioned above, if the wave packets on both sides of the interferometer arm cannot precisely overlap (both in translational and rotational degrees of freedom), they cannot generate interference fringes and thus lead to the loss of contrast, see Japha and Folman (2022) where the analysis of HD effect has been analysed quite nicely including the impact of angular momentum and the libration mode of the NV spin. For macroscopic objects, there is also a challenge of overlap due to the internal degrees of freedom, such as phonon vibration. In a very interesting workHenkel and Folman (2022), the contrast reduction of phonon wave packets induced by coupling NV centres with magnetic fields was demonstrated. More recently, this coupling has been extended to white noise, providing a boundary on creating larger-scale spatial superposition states that have been considered Henkel and Folman (2023).
This work will also investigate the contrast loss caused by phonons in matter-wave interferometers. Unlike atomic-scale interferometers, due to the diamagnetic property of a diamond and its interaction with the magnetic field being proportional to its mass, the diamagnetic interaction becomes essential in the matter-wave interferometers with a massive object. For example, the diamagnetic interaction contributes to the trajectories of the diamond’s centre of mass (CoM) wave packets. Therefore, the magnetic field must be carefully designed to guarantee a close loopMarshman et al. (2022); Zhou et al. (2022a, b, 2023). Besides, there has been recently proposed a mass-independent schemeZhou et al. (2023) based on diamagnetic interaction of creating large superposition size. From the perspective of internal degrees of freedom, the diamagnetic repulsion of the diamonds on either side of the interferometer arms are different, leading to the excitation of varying phonon modes and thus resulting in a contrast reduction of internal degrees of freedom. This is fundamentally similar to the spin-magnetic field interaction papers Henkel and Folman (2022, 2023).
In this regard, our paper differs from the Refs.Henkel and Folman (2022, 2023). We validate their results for the spin-magnetic field interaction. However, we also consider the effect of diamagnetic contribution, which brings another layer of complexity that is not present otherwise. In the case of diamagnetic contribution, the phonons are also coupled to the CoM of the diamond. The difference lies in the fact that for diamagnetic interaction, each carbon atom in the atomic chain on both sides of the interferometer arms experiences repulsion in the same direction, while for the spin-magnetic interaction, the chains are plucked at the NV-centre in opposite directions, as illustrated in Fig. 1. Besides the magnetic effects, we will also briefly mention the role of electric dipole interaction with phonons and how it affects the latter in comparing the loss of contrast for the one-loop interferometer.
The structure of this work is arranged as follows. In section II, we introduced the setup, which includes the phonon model and the statistical model of phonon modes at finite temperature, and introduced the Wigner function. In III, we provided the Hamiltonian of the diamond and presented the model for the trajectory of the diamond wavepacket. In IV, we examined the different effects of diamagnetism-phonon coupling and spin-phonon interaction on phonon dynamics and further utilized displacement operators in Section V to describe the phonon’s contrast reduction. In Section VI, we presented numerical results and compared the contrast loss induced by these two interactions at different temperatures. Also, we have considered contrast loss led by other interactions, potential dipole-electric field interaction, in the level of phonon; this part of the discussion can be found in Appendix D and E.
(a) (a)
(b) (b)
@@
Figure 1: Illustrations of different kinds of forces acting on the lattice chain (represented by black solid wave lines) where positions of atom/spin are dotted in black. The wave packets of the centre of mass (CoMs) in the two interferometry arms denoted here by the left (L) and right (R) arms are plotted by orange and blue circles. Panel (a) shows splitting forces $F_{s}^{\text{L,R}}$ acting on the spin with site number $s$ where they have the same magnitude but opposite direction. Panel (b) shows diamagnetic forces $F_{i}^{\text{L,R}}$ acting on every atom, where, for simplicity, we assume it acts on the spin as well.
@@ II Brief review on Phonon
where $K_{ij}$ is a matrix associated with the inter-atomic interaction while $Q_{q}(X_{i})=Q_{q}^{i}$ represents the eigenvector of the matrix. The position of each atom can be written as $X_{i}(t)+x_{i}(t)$, where the capital $X_{i}(t)$ denotes the equilibrium position of the $i^{\text{th}}$ atom at any instant of time, and $x_{i}$ represents its small vibrations. Because the position of the diamond as a whole varies within the SG apparatus, the equilibrium position $X_{i}(t)$ of each atom in the chain becomes a function of time. Notably, the relative distances between these equilibrium positions, $X_{i+1}-X_{i}=\text{const.}$., remain constant. Therefore, $X_{i}(t)$ can effectively be regarded as the trajectory of diamond’s CoM, with their distances relative to the centre particle $X_{\text{c}}$ remaining unchanged. Here, for convenience, we assume that the NV centre (with equilibrium position $X_{\text{s}}$) is embedded at the CoM position, $X_{\text{s}}(t)=X_{\text{c}}(t)$. In the following, we sometimes neglect the time $t$ for convenience.
The small movement $x_{i}(t)$ can be expanded into mode amplitude $u_{q}$, with creation and annihilation operators as,
@@
For simplicity, in our one-dimensional atomic chain model embedded with an NV centre, we assume that all $N$ particles have an identical atomic mass of carbon $m$ and the mass for the diamond is $M$. The $\omega_{q}$ is the eigenfrequency of the chain model with the fundamental tone,
$\omega_{0}={\pi c}/{L}$ (4)
where $c=17.5\times 10^{3}{\rm m/s}$ is the sound speed within a diamond Jishi, Radi A (2013) and $L$ is the length (estimated via $L=(M/\rho)^{-\frac{1}{3}}$, where the density $\rho=3.51\times 10^{3}$ kg/m3 Greenwood and Earnshaw (2012)). For example, for the masses in a range $M\sim 10^{-14}-10^{-18}$ kg the corresponding fundamental tone ranges from $\omega_{0}\sim 10^{10}-10^{11}$ Hz.
At the initial stage, $t=0$, before the diamond enters the SG apparatus and interacts with the magnetic field, one can assume that the phonons are in a thermal equilibrium state at a temperature $T$ Jishi, Radi A (2013),
@@
where $\sigma_{u_{q}}$ and $\sigma_{\dot{u}_{q}}$ are statistical averages. A review of this detail can be found in the appendix B. In the case of thermal equilibrium, the phonons in each mode $q$ form a mixed state. The amplitudes and momenta of the phonons thus take on Gaussian distributions, represented by the Wigner functionCase (2008) as follows,
$W\left(u_{q},\dot{u}_{q}\right)={\cal{N}}\exp\left[-\left(\frac{u_{q}^{2}}{2\sigma_{u_{q}}^{2}}+\frac{\dot{u}_{q}^{2}}{2\sigma^{2}_{\dot{u}_{q}}}\right)\right],$ (7)
where $\cal{N}$ is normalization factor. Eq. (6) gives the normal mode’s characteristic length and momentum. Note that $\sigma_{u_{q}}$ and $\sigma_{\dot{u}_{q}}$ are statistical averages dependent on temperature $T$, see appendix B. Once we let the diamond experience the SG force, the occupation number of the phonons will evolve, and we will estimate that in the following sections.
@@ III Interactions in Stern-Gerlach Apparatus
For the diamond embedded with a single spin $S$ in the NV, we already assume that it has been prepared in a superposition of opposite direction $S^{\text{R}}=1$ (right) and $S^{\text{L}}=-1$ (left), hence the spatial superposition. The Hamiltonian for the diamond’s CoM consists of three parts  [^2] There is also a term related to $D(S\cdot\hat{n})^{2}$ related to the zero-point splitting Japha and Folman (2022); Zhou et al. (2022a). This term is relevant for the internal spin degrees of freedom. Here, we are not considering this contribution; we do not consider the rotation of the diamond in our current context. For the rotation of NV centre in diamond and its mechanical coupling with magnetic levitation, see Delord et al. (2020); Perdriat et al. (2021).,
@@
the first term is just the kinetic energy relating to CoMs’s momentum. The second term is the the coupling between the NV centre and the magnetic field with the magnetic moment $\mu=g\mu_{\text{b}}$, where the Landé g-factor $g\approx 2$ and Bohr magneton $\mu_{\text{b}}=9.27\times 10^{-24}$ (J/T) O’Handley (1999). For simplicity, we assume the NV is located at the CoM with equilibrium position $X_{\text{s}}=X_{\text{c}}$. The appearance of the third term is because diamond is a diamagnetic material, and its susceptibility is a negative value $\chi_{\rho}=-6.2\times 10^{-9}{\rm m^{3}/kg}$ Heremans et al. (1994) with vacuum permeability $\mu_{0}$. One should note that this term is a macroscopic expression. When a carbon atom chain is placed in an external magnetic field pointing in the $\mathbf{e}_{x}$ direction, each carbon atom will contribute a different magnitude of repulsion according to its equilibrium position $X_{i}$ pointing in the $-\mathbf{e}_{x}$ direction. From the aspect of macroscopic diamond, these small repulsion collectively form a macroscopic diamagnetic force effectively acting on the CoM, $X_{\text{c}}$.
For the magnetic field in Eq. (8), we give its form that is experienced by the $i^{\text{th}}$ atom as,
@@
where $B_{0}$ is the bias magnetic field, and $b(t)$ denotes the gradient in the SG. Note that in Eq.(9), the coordinates along the $\mathbf{e}_{x}$ axis are represented by the uppercase $X_{i}$, indicating the equilibrium positions of atoms in the chain. For example, the magnetic field experienced by the spin, which we have assumed to be located at the centre of the chain, becomes $B_{\text{c}}=B_{0}+b(t)X_{\text{c}}$. In this model, we only consider the one-dimensional case, where the separation of the diamond and the force on the atomic chain is confined to the $\mathbf{e}_{x}$ direction.
We do not consider the motion of the diamond in other dimensions, such as in drop tower experiments where the diamond needs to move in the direction of Earth’s acceleration or in magnetic levitation setups where it moves along a direction orthogonal to $\mathbf{e}_{x}$. In these cases, the form of a magnetic field becomes very complicated for a closed loop, and it needs a careful design, especially when one takes the diamagnetic term into account Zhou et al. (2022c, 2023). For simplicity, we consider the movements of CoMs only along $\mathbf{e}_{x}$, and they are subjected only to the spin-magnetism interaction (second term in Eq. (8)). Therefore, to get a closed loop, the magnetic gradient $b(t)$ can be expressed in a very simple form as Toroš et al. (2021),